7.You buy a bond with $2,000 face value, 1 year to maturity and an 8% coupon rate. You pay $1,900. You hold the bond until maturity. In this case, the yield to maturity is __________ while the rate of return is ___________.

Question options:

8%, 5.25%.

about 13.68%, about 5.26%.

about 13.68%, about 13.68%
Answer #1

as the bond has only 1 yr to maturity

ROR = Yeild to maturity

Paid = 1900 today

Will receive = 2000 + 0.08*2000 = 2160 after 1 yr

using formula P = F /(1+i)^t

1900 = 2160 (1+i)

I = 2160 / 1900 - 1

I = 1.1368 - 1 = 0.1368 = 13.68%

So last option is correct